### v2

# A dialog between cell adhesion and topology at the core of morphogenesis

Abstract
During the development of an organism, cells must coordinate and organize to generate the correct shape, structure, and spatial patterns of tissues and organs, a process known as morphogenesis. The morphogenesis of embryonic tissues is supported by multiple processes that induce the precise physical deformations required for tissues to ultimately form organs with complex geometries. Among the most active players shaping the morphogenetic path are fine-tuned changes in cell adhesion. We review here recent advances showing that changes of a local, pair-wise property defined at the cell-cell contact level has important global consequences for embryonic tissue topology, being determinant in defining both the geometric and material properties of early embryo tissues.
cell adhesion, topology, rigidity, morphogenesis
Introduction
One of the most fundamental constraints underlying the process of morphogenesis is that the acquisition of non-trivial shapes by the forming tissues has to take place without an external agent inducing such changes smith1985developmental ; alberch1989logic ; Wolpert:2015 ; gilbert2023developmental ; sole2024fundamental . In consequence, the different, fine tuned deformations of the embryonic tissue must be the outcome of the collective and coordinated behaviour of different groups of cells which, while changing their mechanical properties in a consistent way, will eventually differentiate and give functionality to the emerging tissue Graner:1993_1 ; gumbiner1996cell ; lecuit2007cell ; oates2009quantitative ; Staple:2010 ; parsons2010cell ; clark2014stresses ; graner2017forms ; tsai2022adhesion . In such processes, underlying core physical properties define the phenotypical frame bi2015density ; Kim:2021 ; atia2021cell ; Petridou:2021 over which selective pressures will act. At the structural level, one must take into account geometrical considerations related to tissue morphology and, even deeper in the identification of the rawest structure, the topological properties of the embryonic tissue Giammona:2021 ; Petridou:2021 ; fischer2023tissues ; Fabreges:2024 . Roughly speaking, the identification of topological properties is a fundamental way to characterize the kind of structure we are dealing with: Whether the structure contains holes, whether the object is tubular-like or spherical-like or, more fundamentally, what the pattern of connection of the different core elements of the system is henle1994combinatorial ; kelley2017general ; Petridou:2021 ; fischer2023tissues . Importantly, the topological characterization is performed regardless of the geometric details, although the topological structure acts as a constraint for the potential geometries. Being topology at such fundamental level of characterization, natural key questions arise: How are the different topological patterns regulated? Is there a feedback between topological structure and material properties? How are different topological phenotypes achieved in front of other potential ones?
Any process producing stress is susceptible to trigger a topological change and other collective phenomena like e.g., active migration bi2016motility ; alert2020physical ; lawson2021jamming or topological sorting graner1992simulation ; Graner:1993_1 ; halbleib2006cadherins ; amack2012knowing ; Maitre:2012 , may as well underlie global tissue structural rearrangements. In this report, we will focus on the role of cell adhesion triggering collective changes on the topology of the tissue organization leading to functional phenotypic traits mcclay1987cell ; takeichi2014dynamic ; Maitre:2015 ; Kim:2021 ; Petridou:2021 ; tsai2022adhesion ; Fabreges:2024 ; Rustarazo:2025 ; Autorino:2025 ; vangheel2026rigidity . We conclude that the dialog established between cell adhesion changes and topology is one of the main drivers of morphogenesis.
Figure 1: Different packings of spheres with their geometry (top) and their associated topologies (bottom). For the sake of simplicity, we only show connected configurations —i.e., with no isolated spheres. With 2 spheres, only one configuration is possible. With 3 spheres, 2 topological configurations are possible, the spheres aligned or arranged in a triangle. Note that the first and the second packings of 3 spheres have different geometries —in the second one a sphere is twisted upwards— but the same topology, because the structure of contacts is the same —a line. With 4 spheres we have 5 different connected topologies in 3D. If the packing topology changes along a given process, either due to external factors or due to changes in the mechanical properties of the spheres, we say that there has been a topological transition.
The topology of the network of cell-cell contacts
One of the most fundamental characterizations of the structure of the embryonic tissue arises from the network of cell-cell contacts escudero2011epithelial ; Petridou:2021 ; fischer2023tissues ; Fabreges:2024 . We will refer to the topology of cell-cell contacts to a particular network structure defined from just considering cells and contacts among them, regardless of any geometric property at any scale. The structure of contacts, when one deals with $n\geq 3$ cells, may have several realizations, since there are several ways to arrange the cells —see fig. 1. Every of such realizations, to which we will refer to as $g_{1},...,g_{N}$, will define a topology of interactions —i.e., a network— whose particular structure is grasped by the adjacency matrix, $A(g_{k})$ newman2018networks . Elements of the adjacency matrix $a_{ij}$ will be proportional to the surface of the contact, in case the cells $c_{i},c_{j}$ are in contact, or zero, in case cells $c_{i},c_{j}$ are not in contact. If we are only interested on the raw topological structure, the elements of the adjacency matrix will be just either one or zero, depending on whether the contact exists or not, respectively. Topological changes, also known as topological transitions —see Box 1—, between structural cell configurations will be represented by changes in the adjacency matrix.
Cell adhesion
Cell adhesion is a fundamental mechanism for maintaining and shaping tissue architecture. Far from acting as a simple mechanical glue, cell adhesion works as a highly regulated mechanochemical process that coordinates tissue structure and maintenance mcclay1987cell ; takeichi1988cadherins ; gumbiner1996cell ; hynes1999cell ; halbleib2006cadherins ; lecuit2007cell ; clark2014stresses ; Winklbauer:2015 ; tsai2022adhesion . Across metazoans, a substantial fraction of the genome encodes proteins associated with cell-cell and cell-matrix adhesion clark1995integrins ; hynes1999cell ; hynes2000evolution . These cell adhesion molecules (CAMs) are typically grouped into several major families, including cadherins, integrins, immunoglobulin superfamily CAMs, and selectins. Generally, CAMs are transmembrane proteins that mediate adhesion outside the cell and connect to the cytoskeleton inside the cell edelman1983cell ; albelda1990integrins ; gumbiner1996cell ; joseph1998cell . Through these connections, adhesion complexes couple external chemical signals, mechanical forces, and tissue-scale tensions to intracellular responses pasternak1989capping ; clark1995integrins ; parsons2010cell ; niessen2011tissue ; salvi2018mechanisms ; de2022interplay ; Campas:2024 .
In spite of the enormous complexity of the adhesion machinery described above, several mechanical phenomena can be studied from a simplified version of the membrane structure —see Fig. 2A. Indeed, since more than a decade, several features of the adhesion dynamics have been studied from the perspective of the theory of foams, establishing an analogy between soap bubbles with potentially variable surface tension properties and cells Weaire:1999 ; Hayashi:2004 ; Maitre:2012 ; Winklbauer:2015 ; Maitre:2015 ; parent2017mechanics ; Kim:2021 ; Fabreges:2024 ; Rustarazo:2025 ; vangheel2026rigidity . In this framework, the key mechanical variable is the relative strength of the cell-fluid $\gamma_{cf}$ against the cell-cell $\gamma_{cc}$ surface tensions, a non-dimensional parameter, $\alpha$, defined as:
$\alpha=\gamma_{cc}/(2\gamma_{cf})\quad.$ (1)
In force-balance conditions the parameter $\alpha$ can be inferred empirically from the contact angle between the two (identical) cells $\theta$ and the fluid as:
$\alpha=\cos\left(\theta/2\right)\quad.$ (2)
The above relation is known as the Young-Dupré relation, and it provides an empirically testable observable connecting theory and experiment —see Fig. 2B. The problem has been generalized to an arbitrary number of cells by defining a Hamiltonian function for $c_{1},..,c_{n}$ cells arranged in a specific configuration $g_{k}$ and a given relative surface tension between cells $\alpha$ as:
$E_{\alpha}(g_{k})=\alpha\sum_{i<j\leq n}a_{ij}+\frac{1}{2}\sum_{i\leq n}s_{i}+K\sum_{i\leq n}\left(v_{i}-v_{0}\right)^{2}\quad,$ (3)
where $a_{ij}$ are elements of the adjacency matrix, corresponding to the contact surface between cells $i$ and $j$, $s_{i}$ stands for the area of cell $i$ in contact with fluid and the last term accounts for the (in)compressibility of the cells around a preferred volume $v_{0}$. We will consider cases in which $K\gg\alpha$, that is, we assume almost perfect volume conservation. We observe that, for $n$ cells, and even in force-balance conditions, some topologies may achieve lower or higher values of $E_{\alpha}$ and that the favoured configurations may change upon changes on $\alpha$.
Figure 2: Cell adhesion between cells. A) View of cell adhesion at different scales of description: micro (membrane level), meso (cell–cell level), and macro (tissue level with connectivity map). B) Changes in cell geometry induced by cell adhesion, characterized by the relative surface tension $\alpha$, observed through the geometrical relationship between intercellular angle $\theta$. C) Topological transition in which cells form a new connection at a critical value due to a reduction in surface tension $\alpha$. Top: geometrical representation of the cell configuration and associated changes in intercellular angle $\theta$. Bottom: Topological representation of the same transition, in which a new link is created.
In consequence, topological transitions, —see fig. 2C— in which e.g., new contacts are formed or old contacts are lost, are expected to occur. The above described topological transitions may differ, although not necessarily, from the so-called T1 transitions occurring in vertex models Farhadifar:2007 ; Staple:2010 ; david2014tissue ; bi2015density ; damavandi2025universality . In T1 transitions, a cell redefines its contact neighbourhood, under the assumption of tissue confluency. In the topological transitions studied here, we do not assume confluency of the tissue and, moreover, some of the properties that will be discussed can only be understood in the case of non-confluent tissues.
Topological rearrangement of early embryo packings
We focus our attention on a case where topological transitions have been reported to be particularly relevant: compaction bowman1970cleavage ; Wolpert:2015 ; gilbert2023developmental . Embryo compaction occurs at different developmental points in mammals, generally around the first 3–6 division rounds after fertilization. Along this process, not only the overall geometry of the blastocyst suffers deep changes, but also the topological structure of cell contacts. The relevance of the topology to understand the biophysics of blastomeres has been highlighted since decades ago hillman1972effect ; graham1979formation ; Giammona:2021 ; kuang2022computable ; skinner2023topological ; Fabreges:2024 . BOX 1: Thermodynamics of topological transitions Consider the set of different potential topologies $g_{1},...,g_{N}$ made of $n$ cells and contacts therein, $\Omega(n)$, each with its energy in equilibrium $E_{\alpha}(g_{i})$ given a relative adhesion strength $\alpha$. The probability of observing a certain configuration $g_{i}$ in equilibrium is: $p_{\alpha}(g_{i})=\frac{e^{-\beta G_{\alpha}(g_{i})}}{\sum\limits_{g_{i}\in\Omega(n)}e^{-\beta G_{\alpha}(g_{i})}}\quad,$ (4) where $\beta$ acts as the inverse of the strength of random fluctuations, $G_{\alpha}(g_{i})$ is the (Gibbs) free energy accounting for the contribution of the energy of the packing $g_{i}$. $G_{\alpha}(g_{i})$ is derived from the contributions of the packing internal energy, $E_{\alpha}(g_{i})$, see eq. (3) and the entropy $S(g_{i})$, which, roughly speaking, accounts for the existing symmetries within the packing meng2010free , leading to: $G_{\alpha}(g_{i})=E_{\alpha}(g_{i})-\frac{1}{\beta}S(g_{i})\quad.$ (5) One can then compute the probability of transitioning between two configurations or microstates $g_{i},g_{j}$ at constant $\alpha$ in a quasi-static way: $\frac{p_{\alpha}(g_{i}\rightarrow g_{j})}{p_{\alpha}(g_{i}\leftarrow g_{j})}=e^{-\beta\left(G_{\alpha}(g_{i})-G_{\alpha}(g_{j})\right)}\quad.$ Under smooth changes of $\alpha$, such balance of transition probabilities may change the equilibrium distribution described in eq. (4). Each equilibrium distribution describes a macrostate and, upon changes in $\alpha$, the probability of transitioning between macrostate $\Sigma$ to macrostate $\Sigma^{\prime}$ obeys the following condition: $\frac{p(\Sigma\rightarrow\Sigma^{\prime})}{p(\Sigma\leftarrow\Sigma^{\prime})}=e^{-\beta(F(\Sigma)-F(\Sigma^{\prime}))}\quad,$ where $F(\Sigma)$ and $F(\Sigma^{\prime})$ stand for the (Helmholtz) free energy of macrostates $\Sigma$ and $\Sigma^{\prime}$, respectively. In consequence, upon changes on $\alpha$, some configurations that were highly probable may become improbable but, in turn, some configurations will become more probable. If this happens, one expects topological rearrangements to take place.
In the 16-cell stage, a difference between outer vs. inner cell populations is established. Exterior cells will later form the trophectoderm, i.e., the extra-embryonic tissues, like the placenta, and inner cells will further develop into the organism itself Wolpert:2015 . In consequence, compaction prepares the system at the 8-cell stage such that, in the next (stochastic) division round, a correct inner-outer cell proportion will be achieved. However, this raises questions like: What is the right configuration at 8-cell stage and how can one identify it? Once this structure is identified, how do embryos undergo such a morphological convergence?
Figure 3: Biological implications of topological transitions. A) Temporal evolution of the relative surface tension $\alpha$ during early mouse embryo development (plot adapted from Fabreges:2024 ). Changes in $\alpha$ have a measurable impact on cell compaction, as observed in experimental 3D confocal reconstructions of the embryo, and at the topological level through the formation of new links in the contact network. This topological transition can be understood and simulated (bottom) by taking into account the experimentally observed changes in cell adhesion $\alpha$ (top). Imaging data of real embryos from T. Hiiragi’s lab. (Scale bar, 25 $\mu m$) B) Temporal evolution of the relative surface tension $\alpha$ during zebrafish embryo development (plot adapted from Petridou:2021 ; Rustarazo:2025 ). Variations in $\alpha$ lead to measurable changes in tissue configuration, as observed experimentally using 2D confocal images, and tissue properties as reflected in network rigidity, where the Giant Rigid Cluster (in red) can either span the entire tissue or cover only a small fraction of it, depending on the relative surface tension $\alpha$. These topological rigid-floppy configurations can be understood and simulated (bottom) by taking into account the experimentally observed values in cell adhesion $\alpha$ (top) Rustarazo:2025 . Imaging data of real embryo tissues from N. Petridou’s lab.
Figure 4: Tissue properties driven by cell adhesion. A) Tissue rigidity transition observed through the phase transition of the Giant Rigid Cluster as a function of the relative surface tension $\alpha$ (plot adapted from Rustarazo:2025 ). B) Formation of tricellular junctions in the tissue, defined as the closure of triangular configurations without interstitial fluid, which also undergoes a transition as a function of the relative surface tension $\alpha$, interestingly occurring at the same critical value $\alpha_{c}$ for GRC (plot adapted from Rustarazo:2025 ). C) Quantification of interstitial fluid in tissues as a function of the relative surface tension $\alpha$, revealing two critical points associated with the closure of triangular configurations ($\alpha_{c}$, tricellular junction formation) and with the corresponding to the closure of residual 3-dimensional fluid pockets where interstitial fluid remains trapped ($\alpha^{\prime}_{c}$, 3D pocket closure) (plot adapted from Autorino:2025 ). D) Surface evolver simulations showing the tissue response to applied pressure, comparing configurations with interstitial fluid to those with closed TCJs. Network size: $25\times 25$, average statistics over 20 replicas, $\alpha=0.92$ for tissue simulation with open TCJs (blue) and $\alpha=0.7$ for tissue simulation with closed TCJs (red), shaded region standard deviation. E) Phase diagram of the tissue states as a function of cell density and relative surface tension. Experimental and simulated tissue configurations are shown across the different regions defined by the critical values ($\phi_{c}$, $\alpha_{c}$). The cell-cell contact network is shown in white and the GRC is highlighted in red. Note that relative surface tension, reflecting cell adhesion, is a better indicator of tissue viscosity and rigidity. Imaging data of real embryo tissues from N. Petridou’s lab, figure adapted from Rustarazo:2025 .
In mouse, at the beginning of the 8-cell stage, a high degree of topological heterogeneity is observed. Then, as long as the compaction process unfolds, all the cell packings become topologically rigid pollaczek1927gliederung ; Laman:1970 ; jacobs1995generic —See Box 2 for terminology and concepts on rigidity— and, thus, the astronomic amount of topological configurations possible with 8 spheres are drastically reduced to 13 Arkus:2009 ; arkus2011deriving —see Box 2. Thanks to the Schöenflies classification for the characterization of chrystallographic structures dorset1995modern , we can identify a converging topology at the end of the compaction phase as the D2d, followed in abundance by the Cs(2) topology Arkus:2009 ; arkus2011deriving . Interestingly, embryos moved through a fairly predictable sequence of topological transitions, reducing topological variability over time Fabreges:2024 .
During compaction, the relative surface tension $\alpha$ —see eq. (1)— decreases Maitre:2012 ; Maitre:2015 ; Fabreges:2024 —see Fig. 3A. This relative increase on cell adhesion is triggered actually by a sharp increase on the cell-fluid surface tension $\gamma_{cf}$ in eq. (1)—, mediated by the actomyosin protein cortex Maitre:2015 . This, in turn, alters the energy balances and, under the frame provided by eq. (3), potentially triggers topological transitions towards more energetically favourable packings —see Fig. 3A. Under increasing adhesion, as observed in real embryos, the D2d topology becomes the lowest-energy configuration, and most initial packings converge to D2d in simulations, often via an intermediate visit to Cs(2) topology, as also observed in reality. To provide a counterfactual validation, it was shown that embryos that lack the maternal allele Myh9, required to generate surface tensions, failed to converge to a specific topology Fabreges:2024 . Interestingly, stochastic variability at the initial stages of the 8-cell period was shown to facilitate the convergence to the D2d structures.
Rigidity transitions at tissue scale regulated by cell adhesion changes
The effect of changes on the (microscopic) mechanical properties of the cells has been studied in paying particular attention on how they change the topology of the tissue and, how this feeds back, in return, into the (macroscopic) material properties of the tissue Kim:2021 ; Petridou:2021 ; Rustarazo:2025 . For larger tissues, the amount of configurations becomes quickly astronomical. However, thanks precisely to the large amount of potential structures, a new approach to the study of the system is possible, considering the potential, expected properties of an ensemble of tissues sharing some structural features that could be i.e., the average number of contacts per cell. This statistical physics-like perspective over the topological properties of tissues has been considered in studying the onset of morphogenesis, right after the cleavage period in Zebrafish development Petridou:2021 ; Rustarazo:2025 . In this precise moment, known as Doming phase, the embryonic tissue switches, in a few minutes, from a high viscosity (solid-like) to a low viscosity (fluid-like) state, enabling the smooth deformation of the embryonic tissue, covering the yolk cell kimmel1995stages —see Fig. 3B. The sudden drop in viscosity petridou2019fluidization resembled a phase transition stanley1971phase ; Petridou:2021 . Analysis of the network structure of cell-cell contacts revealed that, at the fluidization point, the average connectivity dropped slightly below the critical point of generic rigidity —see Box 2—, showing that the framework of rigidity percolation jacobs1995generic could account for the observed behaviour —see Box 2. Using the relation between geometry and mechanical properties provided by the Young Dupré relation —see eq. (1)—, as well as large scale tissue simulations, one can establish a critical point in adhesion, $\alpha_{c}=0.866$..., at which a massive amount of links is formed, triggering a rigidity percolation phase transition —see Box 2— thereby rigidifying the tissue Petridou:2021 ; Rustarazo:2025 —see Fig. 2C and Fig. 3B. In consequence, a slight drop/increase on the cell-cell adhesion is projected, at the tissue level properties, in a highly non-linear way —see Fig. 4A.
Analysis of cell density revealed that, in Wild Type embryos, changes in $\alpha$ potentially inducing rigidification (fluidization) matched also jamming (unjamming) values in cell fraction $\phi$ bolton1990rigidity ; o2002random ; van2010jamming —see Box 2. In consequence, one could argue that crossing the critical point in adhesion actually implies crossing the jamming critical point in cell density, $\phi_{c}\approx 0.84$, and that the actual driver of the rigidification/fluidization processes in non-confluent tissues is related to their density properties. However, it is possible to engineer in vivo embryos whose density lays way beyond the jamming threshold $\phi>\phi_{c}$ that behave fluid-like —a phenomenon observed in tumoral tissues ilina2020cell — and, conversely, engineer in vivo embryos whose tissue density lays below the critical fraction, $\phi<\phi_{c}$ but, still, display a solid-like rheological behaviour Rustarazo:2025 . Conversely, the relative surface tension $\alpha$ reveals as the key driver of the tissue material properties, as no tissue with $\alpha>\alpha_{c}$ behaves solid-like, and no tissue with $\alpha<\alpha_{c}$ behaves fluid-like, regardless of the density $\phi$.
Under these conflicting observations, one must closely explore the effects of crossing $\alpha_{c}$. Aside of the massive creation of cell-cell contacts, a crucial change occurring when crossing $\alpha_{c}$ towards low values is the formation of the three-cell-junctions (TJCs) parent2017mechanics ; Kim:2021 ; Rustarazo:2025 ; Autorino:2025 . Thus, in any triangular configuration of cells, above $\alpha_{c}$, we expect to observe a small pocket of fluid whereas, at values below $\alpha_{c}$, such pocket is closed and a TJC is formed, presumably introducing new structural constraints over the system —see Fig. 4B. Still, if the 3D nature of the tissue is considered, pockets of fluid may remain there, which should be totally expelled at a second critical point, $\alpha_{c}^{\prime}=0.71$... in which a 3D tissue in force balance condition should be totally confluent Autorino:2025 —see Fig. 4C. BOX 2: Topological rigidity, basic concepts - Generic (topological) rigidity: A network structure made of nodes and connecting links is considered to be generically rigid if no independent (geometric) movement of the nodes is possible without stretching/compressing a link. Interpreting links as viscoelastic springs, a generically rigid network will have Young modulus larger than zero, meaning that the structure will react against any deformation. For non-generically rigid networks, deformations are possible at no energy cost. In consequence, generically rigid tissues would be expected to display high viscosity, whereas non-generically rigid (floppy) ones would be expected to display a low viscosity or fluid-like response. - The conditions under which a topological structure is generically rigid depends on the geometrical dimension the structure is embedded in. In 2 dimensions we have that, for a network of $n$ nodes: 1) all nodes must be connected to at least $2$ other nodes and 2) there must be, at least $l$ links, in which $l$: $l_{2D}\geq 2n-3\quad.$ In the case of structures embedded in 3 dimensions, the requirements are that 1) all nodes are connected to at least 3 other nodes and 2) the amount of links $l$ satisfies that: $l_{3D}\geq 3n-6\quad.$ - Rigidity percolation: It is a high order phase transition. Upon a smooth increase of the average number of contacts per cell, a sudden emergence of a generically rigid region that spans almost the whole network, the Giant Rigid Cluster (GRC). A small GRC implies that the structure is floppy, and a large GRC implies that the structure, as a whole, is rigid. In 2 dimensions, the critical point in average the average number of contacts per cell triggering the emergence of the GRC is located at $\langle k_{c}\rangle=4$ for large networks. - Jamming: A system —made e.g., of hard disks— is jammed when the constituents arrest each other and no independent movements are possible. At the jamming critical point in density $\phi$ of hard disks, the network of contacts experiences a sudden emergence of a large GRC —see above. The critical point in density that leads to jamming of hard disks is $\phi_{c}\approx 0.84$, meaning that, in a tiling made of disks spread at random over a 2D plane, if the area covered by disks $\phi$ is such that $\phi\geq\phi_{c}$, the system will be jammed and, thus, will react against any attempt of deformation. Therefore, empirical findings point to the hypothesis that TJCs, triggered at $\alpha_{c}$ , have an active role defining the structure and material response of the tissue. A numerical experiment of tissue compression using Surface evolver software Brakke:1992 shows that, in tissues with the same a-priori cell fraction $\phi>\phi_{c}$, stress propagation —in terms of relative increase of internal pressure in cells— displays a much larger scale in tissues below $\alpha_{c}$, in which TJCs are closed, than in ones above $\alpha_{c}$, the ones containing empty spaces in the central part of triangular structures —see Fig 4D.
As a conclusion, one can construct a phase diagram from the coordinates $\alpha,\phi$ that will be formed by four differentiated regions articulated around a double critical point in adhesion and density $(\alpha_{c},\phi_{c})$ Rustarazo:2025 —see Fig. 4E. These four regions, in turn, may refer to primitive versions of different tissue architectures: 1) $\alpha<\alpha_{c}$, $\phi>\phi_{c}$ epithelial-like 2) $\alpha<\alpha_{c}$, $\phi<\phi_{c}$ lumen-like 3) $\alpha>\alpha_{c}$, $\phi>\phi_{c}$ mesenchymal-like (dense) and 4) $\alpha>\alpha_{c}$, $\phi<\phi_{c}$ mesenchymal-like (sparse) Rustarazo:2025 . In summary, this phase diagram grasps both structural and material properties, and defines a discretized space of possibilities on how these properties may look like in real tissues.
Conclusions and outlook
The theoretical analysis based on tissue topology provides abstract tools to bridge the microscopic level of cell-cell interactions and the emerging macroscopic tissue structure and properties. A central question is, in consequence, how developing organisms regulate tissue topology and, with it, its phenotypic outcomes. This report focuses on cell adhesion, highlighting recent studies which focused on how smooth and continuous changes at cell-cell interfaces propagate nonlinearly at the tissue scale, promoting convergence toward specific phenotypes and, in larger systems, defining a discrete set of possible tissue architectures and material properties. Further research efforts should try to connect the topological/mechanical properties of the tissue with the dynamics of cell fate decision Autorino:2025 or pathological scenarios ilina2020cell ; hannezo2022rigidity , thereby linking function and morphogenesis, the two key processes underlying the developmental path.
Acknowledgements.
We thank Dimitri Fábreges, Takashi Hiiragi and Nicoletta I. Petridou for comments and generously sharing imaging pictures of in vivo embyonic tissues. This work was supported by the Weave project ”Tissue material phase transitions and their role in embryo pattern formation” from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, 518354236, PE 3800/1-1) and Österreichischer Wissenschaftsfonds (FWF, Austrian Science Fund, I6533) (E. F and B. C-M). B.C-M. and A. A-T. acknowledge the support of the field of excellence ”Complexity of life in basic research and innovation” of the University of Graz.